OpenStudy (anonymous):

Find the equation in standard form of the hyperbola if its vertices are at (-1, 2) and (3, 2) and the foci are at (-3, 2) and (5, 2)

OpenStudy (anonymous):

I did (-1,2) and (3,2) so I did this formula with it m=(y2-y1)/(x2-x1). So what I did is 2-2/3-9-3)=0/4=0 m1=o Here is the other one that i did too is (-3,2) and (5,2) so I did the same formula m=(y2-y1)/(x2-x1). So what i did is this 2-2/5-(-3)=0/8=0. m2=0